You invest $3,200 in an account that pays an interest rate of 7.25%, compounded continously.
Calculate the balance of your account after 12 years. Round your answer to the nearest hundredth.

Answer by Boreal(15235)

A=Pe^rt , r is rate, t is time in years
A=3200*e^(.0725*12)=3200*e ^0.87
A= $7638.11